OpenStudy (anonymous):

Medal and Fan/testimonial for a few alg. 2 questions. At tennis practice, Tim practices his backhand and his serve at least 2 hours each day. He works less on his backhand than his serve and practices his serve more than hour daily. Which system of inequalities represents Tim’s daily tennis practice if x represents the number of hours spent practicing his backhand and y represents the number of hours practicing his serve?

OpenStudy (anonymous):

@bibby Bibs.

OpenStudy (bibby):

can I see the answer choices?

OpenStudy (anonymous):

yep one sec.

OpenStudy (anonymous):

They're very long. Hang on. A is x+y>2 x<y y>=1/2

OpenStudy (anonymous):

B is x+y>=2 x>y y>1/2

OpenStudy (bibby):

well for a start let's separate the info \(x+y\ge2\) At tennis practice, Tim practices his backhand and his serve at least 2 hours each day.

OpenStudy (anonymous):

okay.

OpenStudy (bibby):

He works less on his backhand than his serve x<y and practices his serve more than hour daily. y>1

OpenStudy (bibby):

unless that meant 1/2 in which case y>1/2

OpenStudy (anonymous):

That makes D my answer.. Thanks bibs. Another?

OpenStudy (bibby):

or my name's not jeff

OpenStudy (anonymous):

-.- hang on.

OpenStudy (anonymous):

Every year in Delaware there is a contest where people create cannons and catapults designed to launch pumpkins as far in the air as possible. The equation y = 15 + 110x – 16x2 can be used to represent the height, y, of a launched pumpkin, where x is the time in seconds that the pumpkin has been in the air. What is the maximum height that the pumpkin reaches? How many seconds have passed when the pumpkin hits the ground? (Hint: If the pumpkin hits the ground, its height is 0 feet.)

OpenStudy (anonymous):

Okay.. for this one I know to find the seconds I must make the y=0. But I'm stuck on the max height and actually solving it out.

OpenStudy (bibby):

I d on't really remember these. wolframalpha says the maximum is at x=55/16 I assume you'd use the quadratic formula

OpenStudy (anonymous):

thank you. The next is easy, but I'm brain dead. 3+4i/2+4i and you're simplifying.

OpenStudy (bibby):

multiply by the conjugate of the denominator

OpenStudy (anonymous):

Help. Just set it up?

OpenStudy (bibby):

the conjugate of 3+3i is 3-3i

OpenStudy (anonymous):

So multiply (3+4i/2+4i)2-4i

OpenStudy (bibby):

\(\huge \frac{3+4i}{2+4i }*\frac{2-4i}{2-4i}\) we have to multiply both top and bottom by the same thing so that we aren't changing the value of the fraction. when we multiply by x/x it's in essence multiplying by 1

OpenStudy (anonymous):

So how do i do this /.?

OpenStudy (bibby):

just multiply it term by term or use foil if you have to (a+b)(c-d)=(a*c)-(a*d)+(b*c)-(b*d)

OpenStudy (anonymous):

6-16i/4-16i

OpenStudy (bibby):

let's see

OpenStudy (anonymous):

I'm so done with this bibs v.v

OpenStudy (bibby):